Erin's age is 3 times Warren's. In 4 years she will be twice as old as he will be. How old is each? I know Erin is 12 and Warren is 4. How can I check this
Answer by checkley77(12844)

E=3W
E+4=2(W+4)
E+4=2W+8
3W+4=2W+8
3W-2W=8-4
W=4 ANS. FOR WARREN'S AGE.
E=3*4=12 ANS. FOR ERIN'S AGE.
pROOF:
12+4=2(4+4)
16=2*8
16=16